OpenStudy (anonymous):

Show by means of an example that lim x-> a [f(x) + g(x)] may exist even though neither lim x-> a f(x) and lim x -> a g(x) exists.

myininaya (myininaya):

hey

myininaya (myininaya):

The easiest example I can think of is a fraction that equals a constant except at a particular place. And then separate the fraction.

myininaya (myininaya):

Example give me a fraction that doesn't exist at x=1 but exist everywhere else that is a constant value.

myininaya (myininaya):

You can choose that constant value to be 1 if you want...

myininaya (myininaya):

choose a fraction that equals 1 everywhere except at one value of x...

myininaya (myininaya):

oh maybe you aren't here

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \lim _{x\to {\tiny}3}\left[\frac{x}{x-3}+\frac{-3}{x-3}\right] }\) Notice that both of the limits diverge (to \(\small \pm\infty\)) if you take these limits individually/separately. HOWEVER, THE LIMIT OF THE SUM EXISTS!\(\tiny \\[0.8em]\) \(\color{#000000 }{ \displaystyle \lim _{x\to {\tiny}3}\left[\frac{x}{x-3}+\frac{-3}{x-3}\right] =\lim _{x\to {\tiny}3}\left[\frac{x-3}{x-3}\right] =1}\)